OpenStudy (anonymous):

Which of the following statements is true about the graphs of the equations y = (1/2)x − 4 and y = (1.2)x + 10? A. They are perpendicular. B. They intersect at the point (2, −1). C. They intersect at the origin. D. They are parallel.

OpenStudy (anonymous):

1/2* not 1.2

OpenStudy (mathstudent55):

I see your correction. Are these the correct equations? \( y = \dfrac{1}{2}x - 4\) \(y = \dfrac{1}{2} x + 10\)

OpenStudy (anonymous):

yes

OpenStudy (mathstudent55):